OpenStudy (anonymous):

Find the sum of the infinte geometric series: 24 + 12 + 6 + 3 + ... Answer

OpenStudy (anonymous):

The formula for the sum of an infinite geometric series is: \[\frac{a}{1-r}\]where a is the first term, and r is the common ratio. In your problem, a is 24, and r is 1/2 (because 24*(1/2) = 12, 12*(1/2) = 6, etc)

OpenStudy (anonymous):

so is your answer 48

OpenStudy (anonymous):

that is correct :)

OpenStudy (anonymous):

great thanks

OpenStudy (anonymous):

do you know what the answer to this problems is